OpenStudy (anonymous):

How do you solve the following, please show work. If sin(theta)=0.3, find sin (theta+pi)

OpenStudy (freckles):

\[\sin(\theta+\pi)=\sin(\theta) \cos(\pi)+\sin(\pi) \cos(\theta)=?\]

OpenStudy (freckles):

\[\text{So I bet you you are having troubles with the } \cos(\theta) \text{ part } \] Am I right?

OpenStudy (freckles):

We are given \[\sin(\theta)=\frac{3}{10}\] Draw a right triangle giving measurement 10 to the hyp and measurement 3 to the opp

OpenStudy (freckles):

Use the Pythagorean thm to find the adj side :)

OpenStudy (freckles):

\[\text{Then boom you can find } \cos(\theta)\]

OpenStudy (anonymous):

looking at the graph of sin(x) (or using the expansion of sin(a + b)) you can see that sin(x + pi) = -sin(x), just another way to approach it